# AppellateGen: A Benchmark for Appellate Legal Judgment Generation

### v1

@@ Abstract
Hongkun Yang1,2thanks: Equal contribution. Work done while Hongkun Yang was an intern at Nanyang Technological University., Lionel Z. Wang1,3thanks: Equal contribution. Project Leader., Wei Fan4, Yiran Hu5, Lixu Wang1, Chenyu Liu3, Shenghong Fu3, Haoyang Li3, Xin Xu3, Jiexin Zheng6, Wei Dong1thanks: Corresponding author. 1Nanyang Technological University, 2Ocean University of China, 3The Hong Kong Polytechnic University, 4Hong Kong University of Science and Technology, 5Tsinghua University, 6Peking University
@@ 1 Introduction
The integration of Large Language Models (LLMs) into the legal domain has advanced legal NLP, moving from simple clause retrieval to complex reasoning tasks such as passing the Bar Exam (katz2024gpt), and transitioning from discriminative Legal Judgment Prediction (LJP) (cui2023chatlaw; feng2022legal) to Legal Document Generation (li2025casegen; yue2023disc). Despite these advancements, contemporary methodologies predominantly adopt a First-Instance Perspective (zhong2018legal; liu2023ml; xu2020distinguish), where models synthesize judgments directly from static factual descriptions. However, this linear approach neglects a critical phase: the Appellate (Second-Instance) Review (illustrated in Appendix A). Distinct from the fact-to-verdict mapping of a first trial, appellate proceedings are inherently dialectical, necessitating a comparative analysis of the original judgment against appellant grievances to rectify errors in fact-finding or legal application (wechsler1977appellate).
This distinction is particularly pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 SPC2021; SPC2022; SPC2023. This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication, which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
Despite its importance, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench. While recent initiatives have integrated generative tasks li-etal-2025-legalagentbench; li2024lexeval, they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks pipitone2024legalbench and document generation datasets li2025casegen generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
@@
  • Pioneering Task Definition: We extend the frontier of legal NLP from static first-instance judgment prediction to Second-Instance Legal Judgment Generation. This task necessitates a paradigm shift towards dynamic dialectical reasoning, requiring models to identify judicial errors and synthesize conflicting narratives between the original verdict and new evidence.
@@
  • Methodological Framework & Insights: We propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS). By decomposing the generation process into dispute identification, retrieval, prediction, and drafting, our framework significantly mitigates logical hallucinations. Furthermore, our extensive evaluation reveals that general-purpose LLMs with strong reasoning capabilities outperform traditional domain-specific models on this complex task, offering new directions for legal AI development.

### v2

@@ Abstract
Hongkun Yang1,2thanks: Equal contribution.thanks: Work done while Hongkun Yang was an intern at Nanyang Technological University., Lionel Z. Wang1,311footnotemark: 1thanks: Project Leader., Wei Fan4, Yiran Hu5, Lixu Wang1, Chenyu Liu3, Shenghong Fu3, Haoyang Li3, Xin Xu3, Jiexin Zheng6, Wei Dong1thanks: Corresponding author. 1Nanyang Technological University, 2Ocean University of China, 3The Hong Kong Polytechnic University, 4Hong Kong University of Science and Technology, 5Tsinghua University, 6Peking University
@@ 1 Introduction
The integration of Large Language Models (LLMs) into the legal domain has advanced legal NLP, moving from simple clause retrieval to complex reasoning tasks such as passing the Bar Exam (katz2024gpt), and transitioning from discriminative Legal Judgment Prediction (LJP) (cui2023chatlaw; feng2022legal) to Legal Document Generation (li2025casegen; yue2023disc). Despite these advancements, contemporary methodologies predominantly adopt a First-Instance Perspective (zhong2018legal; liu2023ml; xu2020distinguish), where generative models synthesize judgments directly from static factual descriptions. However, this linear approach neglects a critical phase: the Appellate (Second-Instance) Review (illustrated in Appendix LABEL:app:_appellate). Distinct from the fact-to-verdict mapping of a first trial, appellate proceedings are inherently dialectical, necessitating a comparative analysis of the original judgment against appellant grievances to rectify errors in fact-finding or legal application (wechsler1977appellate).
This distinction is particularly pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 SPC2021; SPC2022; SPC2023. This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication (luneburg1981specially; bruhl2010deciding), which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
Despite the importance of appellate review, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench. While recent initiatives have integrated generative tasks li-etal-2025-legalagentbench; li2024lexeval, they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks pipitone2024legalbench and document generation datasets li2025casegen generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
@@
  • Pioneering Task Definition: We extend the frontier of legal NLP from static first-instance judgment prediction to Second-Instance Legal Judgment Generation. This task necessitates a paradigm shift towards dynamic dialectical reasoning, requiring generative models to identify judicial errors and synthesize conflicting narratives between the original verdict and new evidence.
@@
  • Methodological Framework & Insights: We propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS). By decomposing the generation process into dispute identification, retrieval, prediction, and drafting, SLMAS has more outstanding generation performance and mitigates logical hallucinations. Furthermore, our extensive evaluation reveals that general-purpose LLMs with strong reasoning capabilities outperform traditional legal domain-specific models on this complex task.